OpenStudy (anonymous):

Solve the given system of equations. Check your solution. y = 1/2x + 4 2y + 2x = 2

OpenStudy (anonymous):

you can substitute the first equation directly into the second so 2(1/2x+4) +2x =2 distribute the 2 so x+8+2x=2 and solve for x from there

OpenStudy (anonymous):

then once you have a value for x you can put it into either equation to solve for y

OpenStudy (anonymous):

so is it 3,2

OpenStudy (anonymous):

to check if your values are right put the x and y values you obtained into either equation and if they are right then the equation will be true continuing the equation above x+2x=2-8 3x=-6

OpenStudy (anonymous):

so x =-2

OpenStudy (anonymous):

so -2,3

OpenStudy (anonymous):

can you find y from there

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

yes :) and see if you plug x=-2 and y=3 into either equation it proves true, which is how you check your answer

OpenStudy (anonymous):

thanks